OpenStudy (anonymous):

dy/dx = -2sinx + 2cos2x How do I find its critical points?

myininaya (myininaya):

set =0

OpenStudy (anonymous):

Yes I know this step, its just that I get a cos on one side and a sin on the other side. I don't know how to switch them to the same trig function to solve for x

OpenStudy (anonymous):

double angle cos(2x) = 1 - 2sin(x)^2

OpenStudy (anonymous):

ok so the I did the double angle and that gives me: 2sinx = 2(cos^2x - sin^2x) How did you get 1 -2sin(x)^2?

OpenStudy (anonymous):

cos(2x) = 1 - 2sin(x)^2 it's a rule you should have 2sinx = 2 - 2sin^2x

OpenStudy (anonymous):

it's called the double angle formula

OpenStudy (anonymous):

Funny, my book says: Cos2A = cos^2A - sin^2A